# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
The quantum valley Hall effect (QVHE) is characterized by the valley Chern number (VCN) in a way that one-dimensional (1D) chiral metallic states are guaranteed to appear at the domain walls (DW) between two domains with opposite VCN for a given valley. Although in the case of QVHE, the total BC of the system is zero, the BC distributed locally around each valley makes the VCN well-defined as long as inter-valley scattering is negligible. Here, we propose a new type of valley-dependent topological phenomenon that occurs when the BC is strictly zero at each momentum. Such zero Berry curvature (ZBC) QVHE is characterized by the valley Euler number (VEN) which is computed by integrating the Euler curvature around a given valley in two-dimensional (2D) systems with space-time inversion symmetry. 1D helical metallic states can be topologically protected at the DW between two domains with the opposite VENs when the DW configuration preserves either the mirror symmetry with respect to the DW or the combination of the DW space-time inversion , and chiral symmetries. We establish the fundamental origin of ZBC-QVHE. Also, by combining tight-binding model study and first-principles calculations, we propose stacked hexagonal bilayer lattices including h-BX (X=As, P) and large-angle twisted bilayer graphenes as candidate systems with robust helical DW states protected by VEN.
@@
Introduction.— Quantum valley Hall effect (QVHE) is a topological phenomenon arising from nonzero Berry curvature (BC) around two valleys in time-reversal $\mathcal{T}$ symmetric systems with broken inversion $\mathcal{P}$ symmetry [1]. The valley Chern number (VCN) is defined as the integration of Berry curvature (BC) for a given valley, and its change over a domain counts the number of one-dimensional (1D) chiral metallic states at the domain wall (DW) [2, 3, 4]. Although the total BC of the system is zero due to $\mathcal{T}$-symmetry, if the BC is well-localized near each valley and the intervally scattering is negligible, QVHE is robust. Bernal-stacked bilayer graphene is a representative system that exhibits QVHE under vertical electric field breaking $\mathcal{P}$ symmetry [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Many interesting ideas have been proposed to realize QVHE not only in solid state systems [18, 19, 20] but also in classical wave systems [21, 22, 23, 24, 25, 26, 26]. These QVHE setups have potential application for valleytronics [27, 28].
@@
The BC is an essential ingredient in various topological phenomena including the QVHE [29, 30, 31]. However, the recent studies of symmetry-protected topological states have uncovered various intriguing topological states that exist even in the absence of the BC [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. One notable example is the topological states in two-dimensional (2D) systems with space-time inversion ($\mathcal{I_{\text{ST}}}$) symmetry, which appears in the form of either the combination of $\mathcal{T}$ and $\mathcal{P}$ in the absence of spin-orbit coupling or the combination of $\mathcal{T}$ and two-fold rotation about the $z$-axis normal to the 2D plane ($\mathcal{C}_{2z}$), irrespective of the presence or absence of spin-orbit coupling [44]. Although $\mathcal{I_{\text{ST}}}$ symmetry forces the BC to vanish at every momentum, two isolated bands in 2D $\mathcal{I_{\text{ST}}}$ symmetric systems can be characterized by an integer Euler number $\chi$ [45], which induces various intriguing physical properties [46]. For instance, two bands with an Euler number $\chi$ always possess $2\chi$ number of Dirac points between them [45]. Moreover, a topological phase transition changing the Euler number accompanies pair-creation and pair-annihilation of nodal points [45] which can be characterized by non-Abelian braiding processes [47, 48, 46, 49, 50, 51, 52, 53, 54]. Also, the Euler number is a fragile topological index characterizing the band topology of nearly flat bands in twisted bilayer graphene [45, 55, 56, 57], which reduces to a $Z_{2}$ second Stiefel-Whitney number when additional bands are included [44]. However, contrary to the case of stable topological states, the topological Euler bands cannot host boundary in-gap modes unless additional symmetries are imposed to the system [58], as is common in crystalline topological states [59].
@@
Fig. 1(b) describes the band structure evolution of $H_{\eta}$ around a valley. When $m=t_{\perp}=0$, two degenerate Dirac cones, one from the top and the other from the bottom layers, appear at the Fermi energy ($E_{F}=0$) [see Fig. 1(b1)]. Turning on nonzero $m$ generates oppositely signed masses at two Dirac cones, which results in opposite Berry curvature between them [see Fig. 1(b2)]. Nonzero $t_{\perp}$ couples two gapped Dirac cones, which makes the BC vanishing at every momentum [see Fig. 1(b3, b4)]. However, both the upper and lower bands carry a half-integral Euler number, leading to ZBC-QVHE. Note that the presence of a single Dirac point in both the upper and lower bands (the blue dots in Fig. 1(b4)) is also consistent with their half-integral Euler number [47].
Topological invariant.— When interlayer coupling is neglected with $t_{\perp}=0$, each layer carries a quantized VCN with the opposite sign between two layers. Explicitly, the valley Chern number $C_{\tau,\eta}$ for the valley $\eta$ in the layer $\tau$ is $C_{\tau,\eta}=\tfrac{1}{2}\text{sgn}(m\eta\tau)$ where $\tau=+1,-1$ indicate the top and bottom layers, respectively. The change of VCN for each layer (and valley) over the domain wall $\Delta C^{v}_{\tau,\eta}\equiv 2C_{\tau,\eta}=\text{sgn}(m\eta\tau)$ supports gapless DW modes between two domains for the given valley. Since the VCNs of two decoupled layers have the opposite signs, the DW of the bilayer shown in Fig. 2(a) hosts two anti-propagating (helical) in-gap states per valley. The existence of these helical in-gap states at the DW is nothing but the manifestation of the QVHE in each layer  [2, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Interestingly, the helical DW modes survive even when $t_{\perp}\neq 0$ and thus the total VCN vanishes. Below, we show that the helical DW states are protected by the valley Euler number (VEN) of the bilayer.
In $\mathcal{I_{\text{ST}}}$-symmetric systems, one can always find a basis such that both the Hamiltonian and wave functions becomes real [32, 45, 46]. In two-dimensions, two real bands |$\ket{u^{1}(\mathbf{k})}$ and |$\ket{u^{2}(\mathbf{k})}$ can have an integer topological invariant, called the Euler number, defined as
@@
where $\text{Ec}(\mathbf{k})=\bra{\nabla u^{1}(\mathbf{k})}\times\ket{\nabla u^{2}(\mathbf{k})}$ is the Euler curvature and $BZ$ indicates the two-dimensional Brillouin zone [45]. We note that the above integral can also be performed on a disk $\mathcal{D}$ which is equal to the patch Euler number if we neglect 1D integration along the boarder of $\mathcal{D}$ [46, 47]. We choose the size of $\mathcal{D}$ large enough to suppress the Euler curvature along the boundary, then the half-integer value of $\chi_{\mathcal{D}}$ can be determined by integrating the Euler curvature, similar to the valley Chern number calculation by integrating the BC around a valley. For the coupled bilayer Hamiltonian in Eq. (1), the valley Euler number for a valley becomes $\chi_{\eta}=\tfrac{1}{2}\text{sgn}(m\eta)$, and the change of VEN over a DW is $\Delta\chi_{\eta}^{v}\equiv 2\chi_{\eta}=\text{sgn}(m\eta)$. The DW between two domains with opposite VENs can host helical DW states when the system satisfies certain symmetry conditions, as discussed below.
Note that by superposing two real bases, one can define the Chern bases |$\ket{\psi^{\pm}(\mathbf{k})}\equiv(\ket{u^{1}(\mathbf{k})}\pm i\ket{u^{2}(\mathbf{k})})/\sqrt{2}$ such that the Chern number of |$\ket{\psi^{+}(\mathbf{k})}$ (|$\ket{\psi^{-}(\mathbf{k})}$) is equal (opposite) to the Euler number of real bases [46, 60]. When two layers are decoupled with $t_{\perp}=0$, the Chern bases are eigenstates of the Hamiltonian, which give the VCN of each layer. On the other hand, when $t_{\perp}\neq 0$, the Chern bases are not eigenstates anymore, and thus only the VEN is well-defined.
The relation between VCN and VEN is similar to that between the Chern number and $Z_{2}$ invariant in quantum spin Hall systems. When two spin channels with the opposite Chern numbers are superposed, the Chern number of each spin channel is ill-defined if the spin-orbit coupling is present [61, 62, 63] while their Chern number difference is well-defined modulo two when time-reversal symmetry exists [64, 65]. Similarly, in our case, the VCN of each layer is ill-defined when interlayer coupling presents, while the VEN is well-defined when $\mathcal{I_{\text{ST}}}$ symmetry exists
Stability of helical DW states.— The helical DW states related to the VEN can be protected when the DWC satisfies either $M_{y}$ symmetry or the combination of chiral symmetry $\Pi$ and DW space-time inversion $\mathcal{I^{\prime}_{\text{ST}}}\equiv\mathcal{C}_{2z}\mathcal{T}$ (see Fig. 2(a)). Using the Hamiltonian in Eq. (1), the DWC can be constructed by taking $m\rightarrow m(y)$, where $\lim_{|y|\rightarrow\infty}m(y)=\text{sign}(y)m_{0}$ and $m_{0}>0$. Accordingly, the DW is located along the $x$-direction. The Hamiltonian for DWC can be obtained from Eq. (1) by substituting $k_{y}\rightarrow-i\partial_{y}$ as
@@
Figure 2: Realization of ZBC-QVHE in AA’ stacked bilayer h-BX (X= As, P, N). (a) Domain wall configuration (DWC) with $M_{y}$ and $C_{2z}$ symmetries. The blue dashed line and red disc represent the $M_{y}:y\rightarrow-y$ mirror and $C_{2z}$ rotation centre respectively. (b,c,d) Bulk band structure (top) and Euler curvature distribution (bottom) for (b) h-BAs, (c) h-BP, and (d) h-BN. $\mathcal{PT}$ symmetry-protected Dirac cones below the $E_{F}$ are indicated by the purple dots at $K$ and $K^{\prime}$ (e,f,g) DW modes (green lines) for h-BAs, h-BP, and h-BN, respectively, where $\pm$ indicate $M_{y}$ mirror eigenvalues. As the bulk band gap and on-site potential change at the DW increase from (b) to (d), the Euler curvature spreads wider and the helical DW state pair gets closer and pair-annihilate.
@@
For convenience, let us neglect interlayer coupling. Since each layer supports QVHE and the VCN of two layers have opposite signs, helical DW states should appear from each valley in the bilayer. The effective Hamiltonian for the helical DW states from one valley can be written as $H^{\text{DW}}_{0}=v^{\prime}k_{x}\tau_{z}$ where $\tau_{z}$ gives a good quantum number when interlayer coupling is neglected. Since the DW is $M_{y}$ invariant, each DW state should also carry a $M_{y}$ eigenvalue. To have the $M_{y}$ eigenvalue and the layer quantum number simultaneously, $M_{y}$ should be represented by either $\tilde{M}_{y}=\tau_{0}$ or $\tilde{M}_{y}=\tau_{z}$ where the tilde symbol indicates the symmetry representation in the space spanned by helical DW states. $\tilde{M}_{y}=\tau_{z}$ prohibits the mass terms in the form of $H^{\text{DW}}_{m}=m_{1}\tau_{x}+m_{2}\tau_{y}$ even when interlayer coupling is turned on.
@@
The representation $\tilde{M}_{y}=\tau_{z}$ in the DW can be confirmed as follows. For simplicity, let us consider $H^{\text{DW}}(y,\tau)=-iv\partial_{y}\sigma_{y}+\tau m\sigma_{z}$ corresponding to Eq. (3) with $k_{x}=0$ and $t_{\perp}=0$, where $\tau=\pm 1$ indicate the upper and lower layers, respectively. Because of the chiral symmetry $\{H^{\text{DW}}(y,\tau),\sigma_{x}\}=0$, if $\psi(y)$ is a zero mode solution satisfying $H^{\text{DW}}(y,\tau)\psi(y)=0$, $\sigma_{x}\psi(y)$ is also a zero mode solution. Since there is only one zero mode per layer, two solutions should be proportional to each other as $\sigma_{x}\psi(y)=\lambda\psi(y)$ where $\lambda$ is a constant. Namely, $\psi(y)$ is an eigenstate of the chiral symmetry operator $\sigma_{x}$. Moreover, the mass term in $H^{\text{DW}}(y,\tau)$ has the opposite signs in two layers, the DW Hamiltonians for two layers are related as $\sigma_{y}H^{\text{DW}}(y,\tau=+1)\sigma_{y}=H^{\text{DW}}(y,\tau=-1)$. Thus, if $\psi(y)$ is a zero mode solution for one layer, $\sigma_{y}\psi(y)$ is the zero mode solution of the other layer. Since $\psi(y)$ and $\sigma_{y}\psi(y)$ have the opposite $\sigma_{x}$ eigenvalues and $\psi(-y)=\psi(y)$, $\tilde{M}_{y}=\tau_{z}$ can be confirmed (see Supplemental Materials (SM)  [^1] See the Supplemental Materials (SM) for details on tight binding, derivation of edge modes, ZBC QVHE in twisted bilayer graphene, ZBC QVHE in canted antiferromagnetic systems, and DFT computation [82, 83, 84]. for additional discussion). This symmetry representation is not affected by interlayer coupling if it is small enough to maintain the band gap. Therefore, the DWC with $M_{y}$ symmetry can support helical DW states originating from VENs. Since the helical states have the opposite $M_{y}$ eigenvalues, their crossing is stable.
@@
The helical DW states can also be symmetry-protected when $\Pi$ and $\mathcal{I}^{\prime}_{\text{ST}}=C_{2z}T$ satisfying $[\Pi,\mathcal{I^{\prime}_{\text{ST}}}]=0$ exist simultaneously. As noted above, $\Pi$ is a combination of the bipartite chiral symmetry and $M_{z}$ mirror. Also, if the DWC has $C_{2z}$ symmetry, $\mathcal{I^{\prime}_{\text{ST}}}=C_{2z}T$ exists and takes the form of $I^{\prime}_{\text{ST}}=\sigma_{x}\mathcal{K}$ in Eq. (3). Since $\mathcal{I^{\prime}_{\text{ST}}}$ changes the sign of $m$ in Eq. (1), two domains related by $\mathcal{I^{\prime}_{\text{ST}}}$ have the opposite VENs. Since $C_{2z}$ does not mix layers, in the DW $\mathcal{I^{\prime}_{\text{ST}}}$ can be represented by ~$\tilde{\mathcal{I}}^{\prime}_{\text{ST}}=K$ and forces $m_{2}=0$ in $H^{\text{DW}}_{m}=m_{1}\tau_{x}+m_{2}\tau_{y}$. The remaining mass $m_{1}\tau_{x}$ can be annihilated by imposing $\tilde{\Pi}=\tau_{x}$ that satisfies $[\tilde{\Pi},\tilde{\mathcal{I}}^{\prime}_{\text{ST}}]=0$ [67, 68, 69].
@@
Hexagonal bilayer materials.— As an example of AA’ stacked hexagonal bilayers, we consider h-BX (B=Boron and X= As, P, N) as shown in Fig. 2. Fig.2(a) describes the DWC with $M_{y}$ and $C_{2z}$ symmetries composed of two domains with opposite VENs. The band structures of h-BX single domains and the corresponding Euler curvature distribution around two valleys are shown in Fig.2(b,c,d) where we also identify the presence of a Dirac point between the two highest occupied bands at $K$ and $K^{\prime}$ (purple dot in Fig.2(b,c,d)), respectively, demonstrating their half-integral valley Euler number [47]. Our DFT simulation (see SM for detail about the computation method) of the DWC confirms the existence of helical DW modes in h-BAs, and hBP as shown in Fig.2(e), and (f), respectively. Consistent with our prediction, the helical DW modes have opposite $M_{y}$ eigenvalues.
@@
The other reason is due to the abrupt change of the on-site potential at the DW, roughly proportional to the sublattice potential, occurring due to the atomic configuration change as shown in Fig. 2 (a). As h-BN with largest sublattice potential feels stronger potential variation at the DW, the helical DW modes from two valleys can be pair-annihilated. We note that as long as the DWC keeps the $M_{y}$ mirror, the crossing of helical DW states from each valley is stable even if such potential variation happens. Interestingly, this on-site potential variation at the DW can be compensated by applying an external electric field to the domain. As shown in Fig. 3(a, b), the helical DW states can be recovered by applying an external electric field, which is obtained by using a tight-binding model relevant to h-BN with an additional out-of-plane electric field as described in detail in the SM.

### v2

@@ Abstract
The quantum valley Hall effect (QVHE) is characterized by the valley Chern number (VCN) in a way that one-dimensional (1D) chiral metallic states are guaranteed to appear at the domain walls (DW) between two domains with opposite VCN for a given valley. Although in the case of QVHE, the total Berry curvature (BC) of the system is zero, the BC distributed locally around each valley makes the VCN well-defined as long as inter-valley scattering is negligible. Here, we propose a new type of valley-dependent topological phenomenon that occurs when the BC is strictly zero at each momentum. Such zero Berry curvature (ZBC) QVHE is characterized by the valley Euler number (VEN) which is computed by integrating the Euler curvature around a given valley in two-dimensional (2D) systems with space-time inversion symmetry. 1D helical metallic states can be topologically protected at the DW between two domains with the opposite VENs when the DW configuration preserves either the mirror symmetry with respect to the DW or the combination of the DW space-time inversion and chiral symmetries. We establish the fundamental origin of ZBC-QVHE. Also, by combining tight-binding model study and first-principles calculations, we propose stacked hexagonal bilayer lattices including h-BX (X=As, P) and large-angle twisted bilayer graphenes as candidate systems with robust helical DW states protected by VEN.
@@
Introduction.— Quantum valley Hall effect (QVHE) is a topological phenomenon arising from nonzero Berry curvature (BC) around two valleys in time-reversal $\mathcal{T}$ symmetric systems with broken inversion $\mathcal{P}$ symmetry Yao et al. (2009). The valley Chern number (VCN) is defined as the integration of Berry curvature (BC) for a given valley, and its change over a domain counts the number of one-dimensional (1D) chiral metallic states at the domain wall (DW) Zhang et al. (2013); Vaezi et al. (2013); Jung et al. (2011). Although the total BC of the system is zero due to $\mathcal{T}$-symmetry, if the BC is well-localized near each valley and the intervalley scattering is negligible, QVHE is robust. Bernal-stacked bilayer graphene is a representative system that exhibits QVHE under vertical electric field breaking $\mathcal{P}$ symmetry Xiao et al. (2007); Liu et al. (2019); Kim et al. (2014); Martin et al. (2008); Semenoff et al. (2008); Qiao et al. (2011); Pan et al. (2015); Wang et al. (2021); Zhou et al. (2021); Ghader (2020); Ezawa (2013); Kim et al. (2020); Zhu et al. (2019). Many interesting ideas have been proposed to realize QVHE not only in solid state systems Ju et al. (2015); Li et al. (2016); Yin et al. (2016) but also in classical wave systems Qian et al. (2018); Lu et al. (2017); Dong et al. (2017); Noh et al. (2018); Li et al. (2022); Lu et al. (2018, 2018). These QVHE setups have potential application for valleytronics Yao et al. (2008); Yu et al. (2020); Mrudul and Dixit (2021); Rana and Dixit (2023).
@@
The BC is an essential ingredient in various topological phenomena including the QVHE Mondal et al. (2023); Chang et al. (2023); Komatsu et al. (2018). However, the recent studies of symmetry-protected topological states have uncovered various intriguing topological states that exist even in the absence of the BC Zhao and Lu (2017); Ahn et al. (2018a); Ahn and Yang (2019); Ünal et al. (2020); Chen et al. (2021); Ezawa (2021); Zhao et al. (2022); Takahashi and Ozawa (2023); Jiang et al. (2022a); Lim et al. (2023); Park et al. (2021); Kim et al. (2017). One notable example is the topological states in two-dimensional (2D) systems with space-time inversion ($\mathcal{I_{\text{ST}}}$) symmetry, which appears in the form of either the combination of $\mathcal{T}$ and $\mathcal{P}$ in the absence of spin-orbit coupling or the combination of $\mathcal{T}$ and two-fold rotation about the $z$-axis normal to the 2D plane ($\mathcal{C}_{2z}$), irrespective of the presence or absence of spin-orbit coupling Ahn et al. (2019a). Although $\mathcal{I_{\text{ST}}}$ symmetry forces the BC to vanish at every momentum, two isolated bands in 2D $\mathcal{I_{\text{ST}}}$ symmetric systems can be characterized by an integer Euler number $\chi$ [see Eq. (2) for its mathematical definition] Ahn et al. (2019b), which induces various intriguing physical properties Bouhon et al. (2020a).
For instance, for a pair of $I_{ST}$-symmetric bands with the Euler number $\chi$, there must be band crossing points in-between in a way that the total vorticity of band crossing points must be equal to $2\chi$ Ahn et al. (2019b). Moreover, a topological phase transition changing the Euler number accompanies pair-creation and pair-annihilation of nodal points Ahn et al. (2019b) which can be characterized by non-Abelian braiding processes Peng et al. (2022a, b); Bouhon et al. (2020a); Qiu et al. (2023); Jiang et al. (2021); Bouhon et al. (2020b); Slager et al. (2022); Jiang et al. (2022b); Guo et al. (2021). Also, the Euler number is a fragile topological index characterizing the band topology of nearly flat bands in twisted bilayer graphene Ahn et al. (2019b); Bouhon et al. (2019); Po et al. (2019); Song et al. (2019), which reduces to a $Z_{2}$ second Stiefel-Whitney number when additional bands are included Ahn et al. (2019a). However, contrary to the case of stable topological states, the topological Euler bands cannot host boundary in-gap modes unless additional symmetries are imposed to the system Ahn et al. (2019b); Wang et al. (2020), as is common in crystalline topological states Ando and Fu (2015).
@@
Fig. 1(b) describes the band structure evolution of $H_{\eta}$ around a valley. When $m=t_{\perp}=0$, two degenerate Dirac cones, one from the top and the other from the bottom layers, appear at the Fermi energy ($E_{F}=0$) [see Fig. 1(b1)]. Turning on nonzero $m$ generates oppositely signed masses at two Dirac cones, which results in opposite Berry curvature between them [see Fig. 1(b2)]. Nonzero $t_{\perp}$ couples two gapped Dirac cones, which makes the BC vanishing at every momentum [see Fig. 1(b3, b4)]. However, both the upper and lower bands carry a half-integral Euler number, leading to ZBC-QVHE. Note that the presence of a single Dirac point in both the upper and lower bands (the blue dots in Fig. 1(b4)) is also consistent with their half-integral Euler number Peng et al. (2022a).
Topological invariant.— When interlayer coupling is neglected with $t_{\perp}=0$, each layer carries a quantized VCN with the opposite sign between two layers. Explicitly, the valley Chern number $C_{\tau,\eta}$ for the valley $\eta$ in the layer $\tau$ is $C_{\tau,\eta}=\tfrac{1}{2}\text{sgn}(m\eta\tau)$ where $\tau=+1,-1$ indicate the top and bottom layers, respectively. The change of VCN for each layer (and valley) over the domain wall $\Delta C^{v}_{\tau,\eta}\equiv 2C_{\tau,\eta}=\text{sgn}(m\eta\tau)$ supports gapless DW modes between two domains for the given valley. Since the VCNs of two decoupled layers have the opposite signs, the DW of the bilayer shown in Fig. 2(a) hosts two anti-propagating (helical) in-gap states per valley. The existence of these helical in-gap states at the DW is nothing but the manifestation of the QVHE in each layer  Zhang et al. (2013); Kim et al. (2014); Martin et al. (2008); Semenoff et al. (2008); Qiao et al. (2011); Pan et al. (2015); Wang et al. (2021); Zhou et al. (2021); Ghader (2020); Ezawa (2013); Kim et al. (2020). Interestingly, the helical DW modes survive even when $t_{\perp}\neq 0$ and thus the total VCN vanishes. Below, we show that the helical DW states are protected by the valley Euler number (VEN) of the bilayer.
In $\mathcal{I_{\text{ST}}}$-symmetric systems, one can always find a basis such that both the Hamiltonian and wave functions becomes real Zhao and Lu (2017); Ahn et al. (2019b); Bouhon et al. (2020a). In two dimensions, the non-abelian Berry curvature of two real bands, |$\ket{u^{1}(\mathbf{k})}$ and |$\ket{u^{2}(\mathbf{k})}$, forms a real antisymmetric matrix Zhao and Lu (2017); Ahn et al. (2019b). Consequently, the usual Berry curvature, which is the trace of the non-abelian Berry curvature, vanishes, while its Pfaffian remains well-defined and gives the Euler curvature defined as $\text{Ec}(\mathbf{k})=\bra{\nabla u^{1}(\mathbf{k})}\times\ket{\nabla u^{2}(\mathbf{k})}$. By integrating the Euler curvature over the two-dimensional Brillouin zone (BZ), we obtain an integer topological invariant called the Euler number, given by
@@
We note that the above integral can also be performed on a disk $\mathcal{D}$ which is equal to the patch Euler number if we neglect 1D integration along the boarder of $\mathcal{D}$ Bouhon et al. (2020a); Peng et al. (2022a). We choose the size of $\mathcal{D}$ large enough to suppress the Euler curvature along the boundary, then the half-integer value of $\chi_{\mathcal{D}}$ can be determined by integrating the Euler curvature, similar to the valley Chern number calculation by integrating the BC around a valley. For the coupled bilayer Hamiltonian in Eq. (1), the valley Euler number for a valley becomes $\chi_{\eta}=\tfrac{1}{2}\text{sgn}(m\eta)$, and the change of VEN over a DW is $\Delta\chi_{\eta}^{v}\equiv 2\chi_{\eta}=\text{sgn}(m\eta)$. The DW between two domains with opposite VENs can host helical DW states when the system satisfies certain symmetry conditions, as discussed below.
Note that by superposing two real bases, one can define the Chern bases |$\ket{\psi^{\pm}(\mathbf{k})}\equiv(\ket{u^{1}(\mathbf{k})}\pm i\ket{u^{2}(\mathbf{k})})/\sqrt{2}$ such that the Chern number of |$\ket{\psi^{+}(\mathbf{k})}$ (|$\ket{\psi^{-}(\mathbf{k})}$) is equal (opposite) to the Euler number of real bases Bouhon et al. (2020a); Guan et al. (2022). When two layers are decoupled with $t_{\perp}=0$, the Chern bases are eigenstates of the Hamiltonian, which give the VCN of each layer. On the other hand, when $t_{\perp}\neq 0$, the Chern bases are not eigenstates anymore, and thus only the VEN is well-defined.
The relation between VCN and VEN is similar to that between the Chern number and $Z_{2}$ invariant in quantum spin Hall systems. When two spin channels with the opposite Chern numbers are superposed, the Chern number of each spin channel is ill-defined if the spin-orbit coupling is present Kane and Mele (2005a); Bernevig and Zhang (2006); Sheng et al. (2005) while their Chern number difference is well-defined modulo two when time-reversal symmetry exists Kane and Mele (2005b); Fu and Kane (2007). Similarly, in our case, the VCN of each layer is ill-defined when interlayer coupling presents, while the VEN is well-defined when $\mathcal{I_{\text{ST}}}$ symmetry exists
Stability of helical DW states.— The helical DW states related to the VEN can be protected when the DWC satisfies either $M_{y}$ symmetry or the combination of chiral symmetry $\Pi$ and DW space-time inversion $\mathcal{I^{\prime}_{\text{ST}}}\equiv\mathcal{C}_{2z}\mathcal{T}$ (see Fig. 2(a)). The former ($M_{y}$ symmetry) is easily achievable and enough to protect the DW states while the latter also exists in various model Hamiltonians. Using the Hamiltonian in Eq. (1), the DWC can be constructed by taking $m\rightarrow m(y)$, where $\lim_{|y|\rightarrow\infty}m(y)=\text{sign}(y)m_{0}$ and $m_{0}>0$. Accordingly, the DW is located along the $x$-direction. The Hamiltonian for DWC can be obtained from Eq. (1) by substituting $k_{y}\rightarrow-i\partial_{y}$ as
@@
Figure 2: Realization of ZBC-QVHE in AA’ stacked bilayer h-BX (X= As, P, N). (a) Domain wall configuration (DWC) with $M_{y}$ and $C_{2z}$ symmetries. The blue dashed line and red disc represent the $M_{y}:y\rightarrow-y$ mirror and $C_{2z}$ rotation center, respectively. (b,c,d) Bulk band structure (top) and Euler curvature distribution (bottom) for (b) h-BAs, (c) h-BP, and (d) h-BN, respectively. $\mathcal{PT}$ symmetry-protected Dirac cones below the $E_{F}$ are indicated by the purple dots at $K$ and $K^{\prime}$ (e,f,g) DW modes (green lines) for h-BAs, h-BP, and h-BN, respectively, where $\pm$ indicate $M_{y}$ mirror eigenvalues. As the bulk band gap increases (b to d), the Euler curvature spreads wider, and the crossing points between green lines get closer and eventually pair-annihilated (e to g).
@@
For convenience, let us first neglect interlayer coupling. Since each layer supports QVHE and the VCN of two layers have opposite signs, helical DW states should appear from each valley in the bilayer. The effective Hamiltonian for the helical DW states from one valley can be written as $H^{\text{DW}}_{0}=v^{\prime}k_{x}\tau_{z}$ where $\tau_{z}$ gives a good quantum number when interlayer coupling is neglected. Since the DW is $M_{y}$ invariant, each DW state should also carry a $M_{y}$ eigenvalue. To have the $M_{y}$ eigenvalue and the layer quantum number simultaneously, $M_{y}$ should be represented by either $\tilde{M}_{y}=\tau_{0}$ or $\tilde{M}_{y}=\tau_{z}$ where the tilde symbol indicates the symmetry representation in the space spanned by helical DW states. $\tilde{M}_{y}=\tau_{z}$ prohibits the mass terms in the form of $H^{\text{DW}}_{m}=m_{1}\tau_{x}+m_{2}\tau_{y}$ even when interlayer coupling is turned on.
@@
The representation $\tilde{M}_{y}=\tau_{z}$ in the DW can be confirmed as follows. For simplicity, let us consider $H^{\text{DW}}(y,\tau)=-iv\partial_{y}\sigma_{y}+\tau m\sigma_{z}$ corresponding to Eq. (3) with $k_{x}=0$ and $t_{\perp}=0$, where $\tau=\pm 1$ indicate the upper and lower layers, respectively. Because of the chiral symmetry $\{H^{\text{DW}}(y,\tau),\sigma_{x}\}=0$, if $\psi(y)$ is a zero mode solution satisfying $H^{\text{DW}}(y,\tau)\psi(y)=0$, $\sigma_{x}\psi(y)$ is also a zero mode solution. Since there is only one zero mode per layer, two solutions should be proportional to each other as $\sigma_{x}\psi(y)=\lambda\psi(y)$ where $\lambda$ is a constant. Namely, $\psi(y)$ is an eigenstate of the chiral symmetry operator $\sigma_{x}$. Moreover, as the mass term in $H^{\text{DW}}(y,\tau)$ has the opposite signs in two layers, the DW Hamiltonians for two layers are related as $\sigma_{y}H^{\text{DW}}(y,\tau=+1)\sigma_{y}=H^{\text{DW}}(y,\tau=-1)$. Thus, if $\psi(y)$ is a zero mode solution for one layer, $\sigma_{y}\psi(y)$ is the zero mode solution of the other layer. Since $\psi(y)$ and $\sigma_{y}\psi(y)$ have the opposite $\sigma_{x}$ eigenvalues and $\psi(-y)=\psi(y)$, $\tilde{M}_{y}=\tau_{z}$ can be confirmed (see Supplemental Materials (SM)  [^1] See the Supplemental Materials (SM) for details on tight binding, derivation of edge modes, ZBC QVHE in twisted bilayer graphene, ZBC QVHE in canted antiferromagnetic systems, DW stability study, and DFT computation which includes Refs.  Mondal et al. (2023); Blöchl (1994); Kresse and Hafner (1993); Kresse and Joubert (1999); Mele (2012); Shallcross et al. (2008); Park et al. (2019); Mele (2010); Kindermann (2015); Liu et al. (2022); Ahn et al. (2018b); Schubert et al. (2012); Wu et al. (2013); Queiroz and Schnyder (2015); Pieper and Fehske (2016); Araújo and Andrade (2019); Ghadimi et al. (2021). for additional discussion). This symmetry representation is not affected by interlayer coupling if it is small enough to maintain the band gap. Therefore, the DWC with $M_{y}$ symmetry can support helical DW states originating from VENs. Since the helical states have the opposite $M_{y}$ eigenvalues, their crossing is stable.
@@
The helical DW states can also be symmetry-protected when $\Pi$ and $\mathcal{I}^{\prime}_{\text{ST}}=C_{2z}T$ satisfying $[\Pi,\mathcal{I^{\prime}_{\text{ST}}}]=0$ exist simultaneously. As noted above, $\Pi$ is a combination of the bipartite chiral symmetry and $M_{z}$ mirror. Also, if the DWC has $C_{2z}$ symmetry, $\mathcal{I^{\prime}_{\text{ST}}}=C_{2z}T$ exists and takes the form of $I^{\prime}_{\text{ST}}=\sigma_{x}\mathcal{K}$ in Eq. (3). Since $\mathcal{I^{\prime}_{\text{ST}}}$ changes the sign of $m$ in Eq. (1), two domains related by $\mathcal{I^{\prime}_{\text{ST}}}$ have the opposite VENs. Since $C_{2z}$ does not mix layers, in the DW $\mathcal{I^{\prime}_{\text{ST}}}$ can be represented by ~$\tilde{\mathcal{I}}^{\prime}_{\text{ST}}=K$ and forces $m_{2}=0$ in $H^{\text{DW}}_{m}=m_{1}\tau_{x}+m_{2}\tau_{y}$. The remaining mass $m_{1}\tau_{x}$ can be annihilated by imposing $\tilde{\Pi}=\tau_{x}$ that satisfies $[\tilde{\Pi},\tilde{\mathcal{I}}^{\prime}_{\text{ST}}]=0$ Bzdušek and Sigrist (2017); Kim et al. (2021); Mondal et al. (2022).
@@
Hexagonal bilayer materials.— As an example of AA’ stacked hexagonal bilayers, we consider h-BX (B=Boron and X= As, P, N) as shown in Fig. 2. Fig.2(a) describes the DWC with $M_{y}$ and $C_{2z}$ symmetries composed of two domains with opposite VENs. The band structures of h-BX single domains and the corresponding Euler curvature distribution around two valleys are shown in Fig.2(b,c,d) where we also identify the presence of a Dirac point between the two highest occupied bands at $K$ and $K^{\prime}$ (purple dot in Fig.2(b,c,d)), respectively, demonstrating their half-integral valley Euler number Peng et al. (2022a). Our DFT simulation (see SM for detail about the computation method) of the DWC confirms the existence of helical DW modes in h-BAs, and hBP as shown in Fig.2(e), and (f), respectively. Consistent with our prediction, the helical DW modes have opposite $M_{y}$ eigenvalues.
@@
The other reason is due to the abrupt change of the on-site potential at the DW, roughly proportional to the sublattice potential, occurring due to the atomic configuration change as shown in Fig. 2 (a). As h-BN with the largest sublattice potential feels stronger potential variation at the DW, the helical DW modes from two valleys can be pair-annihilated. We note that as long as the DWC keeps the $M_{y}$ mirror, the crossing of helical DW states from each valley is stable even if such potential variation happens. Interestingly, this on-site potential variation at the DW can be compensated by applying an external electric field to the domain. As shown in Fig. 3(a, b), the helical DW states can be recovered by applying an external electric field, which is obtained by using a tight-binding model relevant to h-BN with an additional out-of-plane electric field as described in detail in the SM.